# Antenna Coding Empowered by Pixel Antennas

### v1

@@ Abstract
Index Terms:
Antenna coding, beamspace, binary optimization, capacity, channel gain, codebook, MIMO, pixel antenna.
@@ I Introduction
Pixel antennas are a flexible antenna technology to design highly reconfigurable antennas and dates back to 2004 [2]. The concept of the pixel antenna is based on discretizing a continuous radiation surface into small elements, which are referred to as pixels, and connecting adjacent pixels through hardwires or RF switches [3], [4], [5], [6] and [7]. By controlling the connections between pixels, various antenna topologies can be formed so that the pixel antenna can be reconfigured to achieve a wide range of distinct antenna characteristics such as operating frequency, radiation pattern, and polarization. For example, frequency-reconfigurable pixel antennas, which can reconfigure the operating frequency from 1.56 GHz to 3.3 GHz and from PCS band to UMTS band, have been designed in [3] and [4], and pattern-reconfigurable pixel antennas, which can reconfigure the radiation pattern to implement $360\text{\textdegree}$ single- or multi-beam steering, have been designed in [5], [6], and [7]. Moreover, various pixel antennas have been designed for different applications. In [8], a reconfigurable intelligent surface with pixelated elements, which can reconfigure the phase shift of reflected waves, has been designed. In [9], [10], and [11], multiport pixel rectennas, which can maximize the output dc power for ambient RF energy harvesting, have been developed. In [12], MIMO antennas with pixelated surface, which provide low mutual coupling and spatial correlation to maximize the ergodic capacity, have been developed. In [13], defected ground structures with pixelated grid, which suppress cross-polarization and achieve circular polarization, have been designed. While these works [3], [4], [5], [6], [7], [8], [9], [10], [11], [12], [13] have designed pixel antennas with significant performance enhancement, the investigations are limited at the level of antenna hardware design and do not consider the impact on wireless communications at the level of the system. Thus, it is necessary to explore using pixel antennas to design and optimize wireless communication systems.
@@
Organization: Section II introduces the pixel antenna model and antenna coding. Section III presents the SISO pixel antenna system and antenna coding design for channel gain maximization. Section IV presents the MIMO pixel antenna system and antenna coding design for capacity maximization. Section V provides performance evaluation for the SISO and MIMO pixel antenna system. Section VI concludes this work.
Notation: Bold lower and upper case letters represent vectors and matrices, respectively. A symbol without bold font denotes a scalar. $\mathbb{R}$ and $\mathbb{C}$ represent real and complex number sets, respectively. $\mathbb{E}\left[\cdotp\right]$ denotes the expectation. $\left(\cdotp\right)^{+}$denotes the positive part. $a^{\ast}$ and $\left|a\right|$ denote the conjugate and modulus of a complex number $a$, respectively. $\left[\mathbf{a}\right]_{i}$ and $\left\|\mathbf{a}\right\|$ denote the $i$th element and $l_{2}$-norm of a vector $\mathbf{a}$, respectively. $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, $\left[\mathbf{A}\right]_{:,i}$, $\left[\mathbf{A}\right]_{i,j}$, $\left|\mathbf{A}\right|$, and $\textrm{Tr}\left(\mathbf{A}\right)$ denote the transpose, conjugate transpose, $i$th column, $\left(i,j\right)$th element, determinant, and trace of a matrix $\mathbf{A}$, respectively. $\mathcal{CN}(\boldsymbol{0},\boldsymbol{\Sigma})$ denotes the circularly symmetric complex Gaussian distribution with mean 0 and covariance matrix $\boldsymbol{\Sigma}$. $\text{diag}\left(a_{1},...,a_{N}\right)$ denotes a diagonal matrix with entries $a_{1},...,a_{N}$. $\mathbf{I}$ denotes an identity matrix.
@@ II Pixel Antenna
The pixel antenna can be modeled using microwave multiport network theory. Specifically, considering a pixel antenna which has $Q$ switches, we can model the pixel antenna as a multiport circuit network by replacing the $Q$ switches with $Q$ ports as shown in Fig. 1(b). Thus, the entire pixel antenna can be modeled as a $(Q+1)$-port circuit network, which consists of one antenna port and $Q$ pixel ports denoting the connections between adjacent pixels (which are numbered 2 to $Q+1$), as shown in Fig. 3. The $(Q+1)$-port circuit network can be characterized by a $(Q+1)\times(Q+1)$ impedance matrix $\mathbf{Z}$, which can be partitioned as
@@
[…]
@@
We group $z_{\textrm{L},q}$ $\forall q$ into a diagonal load impedance matrix $\mathbf{Z}_{\textrm{L}}=\textrm{diag}(z_{\textrm{L},1},\ldots,z_{\textrm{L},Q})\in\mathbb{C}^{Q\times Q}$ and group $b_{q}$ $\forall q$ into a vector $\mathbf{b}=[b_{1},\ldots,b_{Q}]^{T}\in\mathbb{R}^{Q\times 1}$, which represents the states of all switches and is referred to as an antenna coder. Therefore, the load impedance matrix $\mathbf{Z}_{\textrm{L}}$ can be coded by $\mathbf{b}$, denoted as $\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)$, so that $\mathbf{v}_{\textrm{P}}$ and $\mathbf{i}_{\textrm{P}}$ can be related by
@@
The load impedance matrix $\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)$ can be coded by the antenna coder $\mathbf{b}$, which enables coding of the currents $\mathbf{i}\left(\mathbf{b}\right)$ and radiation pattern of the pixel antenna $\mathbf{e}\left(\mathbf{b}\right)$. There are in total $2^{Q}$ different combinations for the antenna coders and each antenna coder produces distinct radiation patterns. As a result, we can select the antenna coder $\mathbf{b}$ among a wide range of $2^{Q}$ combinations to find the optimum radiation pattern for wireless systems. In the following sections, we will introduce SISO and MIMO pixel antenna systems and show the antenna coding design for enhancing wireless systems.
@@ III-A Beamspace Channel Model
The SISO pixel antenna system is equipped with a conventional antenna at the transmitter and a pixel antenna at the receiver, as shown in Fig. 4. To model the channel for the SISO pixel antenna system, we use the beamspace channel representation [29]. Specifically, utilizing a virtual channel representation, we can represent the beamspace channel model in the angular domain, written as
$h\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}$ (8)
@@ III-B Antenna Coding Design
$\left|h\left(\mathbf{b}_{\textrm{R}}\right)\right|^{2}=\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (11)
Assuming the channel state information (CSI) for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ is perfectly known as the CSI can be estimated by using methods in [30], [31], we aim to optimize the antenna coder for the pixel antenna $\mathbf{b}_{\textrm{R}}$ to maximize the channel gain, formulated as
$\underset{\mathbf{b}_{\textrm{R}}}{\text{max}\ \ }$ $\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}$ (12)
@@
Various algorithms have been proposed and utilized for pixel antenna design and optimization such as the genetic algorithm [4], [9], mixed integer linear programming [32], N-port characteristic mode analysis [33], [34], perturbation sensitivity analysis [35], and the adjoint method combined with the method of moving asymptote [36]. In this work, to solve the problem (12)-(13), we use an efficient algorithm called the Successive Exhaustive Boolean Optimization (SEBO) [37], which has two steps: 1) cyclically optimizing each block of the binary variables by exhaustive search until convergence, and 2) randomly flipping bits in the converged binary solution to check if there is any other local optimum with a better objective value. More details on SEBO can be found in [37].
@@ III-C Codebook Design for Antenna Coding
$\mathbf{b}_{\textrm{R}}^{\star}=\underset{\mathbf{b}_{\textrm{R}}\in\mathcal{B}}{\text{argmax}}\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (14)
@@
$\underset{\mathbf{b}_{m}}{\text{max}}$ $\mathbb{E}\left[\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}\right]$ (15)
@@
The codebook design problem (15) and (16) is related to the vector quantization problem [38]. Specifically, we assume a training set containing $L$ realizations for the virtual channel matrix, given by
@@
where $\mathbf{H}_{\textrm{V}}^{\left[l\right]}$ denotes the $l$th virtual channel realization. Leveraging the sample average approximation, we can approximate the expectation of channel gain in (15) as
$\mathbb{E}\left[\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}\right]\approx\frac{1}{L}\sum_{l=1}^{L}\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (18)
@@
$\mathcal{H}_{m}$ =$=\left\{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mid\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}\right.$
$\left.\qquad\hfill\geq\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{m^{\prime}}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall l,\forall m^{\prime}\neq m\right\},\forall m.$ (19)
@@
$\underset{\mathbf{b}_{m}}{\text{max}}$ $\sum_{m=1}^{M}\sum_{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\in\mathcal{H}_{m}}\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}$ (20)

### v2

@@ I Introduction
Pixel antennas are a flexible antenna technology to design highly reconfigurable antennas and dates back to 2004 [2], [3]. The concept of the pixel antenna is based on discretizing a continuous radiation surface into small elements, which are referred to as pixels, and connecting adjacent pixels through hardwires or RF switches [4, 5, 6], [7], [8], [9] and [10]. By controlling the connections between pixels, various antenna topologies can be formed so that the pixel antenna can be reconfigured to achieve a wide range of distinct antenna characteristics such as operating frequency, radiation pattern, and polarization. For example, a frequency-reconfigurable pixel antenna, which can reconfigure the operating frequency from PCS band to UMTS band, has been designed in [7], and pattern-reconfigurable pixel antennas, which can reconfigure the radiation pattern to implement $360\text{\textdegree}$ single- or multi-beam steering, have been designed in [8], [9], and [10]. Moreover, various pixel antennas have been designed for different applications. In [11], a reconfigurable intelligent surface with pixelated elements, which can reconfigure the phase shift of reflected waves, has been designed. In [12], and [13], multiport pixel rectennas, which can maximize the output dc power for ambient RF energy harvesting, have been developed. In [14], MIMO antennas with pixelated surface, which provide low mutual coupling and spatial correlation to maximize the ergodic capacity, have been developed. In [15], defected ground structures with pixelated grid, which suppress cross-polarization and achieve circular polarization, have been designed. While these works [3], [4, 5, 6], [7], [8], [9], [10], [11], [12], [13], [14], [15] have designed pixel antennas with significant performance enhancement, the investigations are limited at the level of antenna hardware design and do not consider the impact on wireless communications at the level of the system. Thus, it is necessary to explore using pixel antennas to design and optimize wireless communication systems.
@@
Organization: Section II introduces the pixel antenna model and antenna coding. Section III presents the SISO pixel antenna system and antenna coding design for channel gain maximization. Section IV presents the MIMO pixel antenna system and antenna coding design for capacity maximization. Section V provides performance evaluation for the SISO and MIMO pixel antenna system. Section VI provides discussions and future work. Section VII concludes this work.
[…]
@@ II Pixel Antenna
The pixel antenna can be modeled using microwave multiport network theory [^1] The multiport circuit network model can be traced back to the method of moments (MoM) [31] which uses an impedance matrix to model an antenna. Using multiport circuit network to model lumped elements, such as capacitor, inductor, resistor, and RF switches, has been developed and demonstrated in [32], [33] and verified in reconfigurable antenna designs [7], [34].. Specifically, considering a pixel antenna which has $Q$ switches, we can model the pixel antenna as a multiport circuit network by replacing the $Q$ switches with $Q$ ports as shown in Fig. 1(b). Thus, the entire pixel antenna can be modeled as a $(Q+1)$-port circuit network, which consists of one antenna port and $Q$ pixel ports denoting the connections between adjacent pixels (which are numbered 2 to $Q+1$), as shown in Fig. 3. The $(Q+1)$-port circuit network can be characterized by a $(Q+1)\times(Q+1)$ impedance matrix $\mathbf{Z}$, which can be partitioned as
@@
The switch connected to the $q$th pixel port has two states, on or off [^2] It is the RF switch that has on or off states and that it is not the pixel element that has on or off states. The RF switch in off state will only make the current at the RF switch zero, but this does not mean that the surface current through the pixel element will disappear, which is because pixel elements are closely coupled with each other. In addition, control of the RF switch states can be achieved by ensuring a dc path exists through the structure by using RF chokes.. As shown in Fig. 3, we can model the switch as a load impedance $z_{\textrm{L},q}$ $\forall q$ with open or short circuit, i.e. $z_{\textrm{L},q}=0$ or $\infty$. Therefore, a binary variable $b_{q}\in\left\{\textrm{0,1}\right\}$ can be used to denote the state of the switch connected to the $q$th port, i.e.
@@
Numerically, for the switch off state, we can set $z_{\textrm{L},q}$ as a very large value, for example $z_{\textrm{L},q}=10^{8}$, to approach the infinite. We group $z_{\textrm{L},q}$ $\forall q$ into a diagonal load impedance matrix $\mathbf{Z}_{\textrm{L}}=\textrm{diag}(z_{\textrm{L},1},\ldots,z_{\textrm{L},Q})\in\mathbb{C}^{Q\times Q}$ and group $b_{q}$ $\forall q$ into a vector $\mathbf{b}=[b_{1},\ldots,b_{Q}]^{T}\in\mathbb{R}^{Q\times 1}$, which represents the states of all switches and is referred to as an antenna coder. Therefore, the load impedance matrix $\mathbf{Z}_{\textrm{L}}$ can be coded by $\mathbf{b}$, denoted as $\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)$, so that $\mathbf{v}_{\textrm{P}}$ and $\mathbf{i}_{\textrm{P}}$ can be related by
@@
The load impedance matrix $\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}\right)$ can be coded by the antenna coder $\mathbf{b}$, which enables coding of the currents $\mathbf{i}\left(\mathbf{b}\right)$ and radiation pattern of the pixel antenna $\mathbf{e}\left(\mathbf{b}\right)$. Thus, given the impedance matrix $\mathbf{Z}$ and open-circuit radiation pattern matrix $\mathbf{E}_{\textrm{oc}}$, we can find the radiation pattern of pixel antenna $\mathbf{e}\left(\mathbf{b}\right)$ configured by any antenna coder by (6) and (7) with low computational complexity of $\mathcal{O}\left(Q^{3}+Q^{2}+Q+2K(Q+1)\right)$. In Section V-D, we will show that this multiport circuit network model has the same accuracy as the full-wave EM simulation but requires significantly less computational time.
In total, there are $2^{Q}$ different combinations for the antenna coders and each antenna coder produces distinct radiation patterns. As a result, we can select the antenna coder $\mathbf{b}$ among a wide range of $2^{Q}$ combinations to find the optimum radiation pattern for wireless systems. In the following sections, we will introduce SISO and MIMO pixel antenna systems and show the antenna coding design for enhancing wireless systems.
@@ III-A Beamspace Channel Model
The SISO pixel antenna system is equipped with a conventional antenna at the transmitter and a pixel antenna at the receiver, as shown in Fig. 4. To model the channel for the SISO pixel antenna system, we use the beamspace channel representation [35], [36]. Specifically, utilizing a virtual channel representation, we can represent the beamspace channel model in the angular domain, written as
$h\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}$ (8)
@@ III-B Antenna Coding Design
$\left|h\left(\mathbf{b}_{\textrm{R}}\right)\right|^{2}=\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (11)
Assuming the channel state information (CSI) for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ is perfectly known as the CSI can be estimated by using methods in [37], we aim to optimize the antenna coder for the pixel antenna $\mathbf{b}_{\textrm{R}}$ to maximize the channel gain, formulated as
$\underset{\mathbf{e}_{\textrm{R}},\mathbf{b}_{\textrm{R}},\mathbf{i}_{\textrm{R}},i_{\textrm{A}}^{\textrm{R}}}{\text{max}}$ $\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}$ (12)
s.t. $\left\|\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)\right\|=1,$ (13)
$\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{E}_{\textrm{oc}}\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right),$ (14)
$\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\left[\begin{array}[]{c}1\\ -\left(\mathbf{Z}_{\textrm{PP}}+\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}_{\textrm{R}}\right)\right)^{-1}\mathbf{z}_{\textrm{PA}}\end{array}\right]i_{\textrm{A}}^{\textrm{R}},$ (17)
$\left[\mathbf{b}_{\textrm{R}}\right]_{q}\in\left\{0,1\right\},\forall q,$ (18)
where $\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ is the currents at all ports of the pixel antenna coded by $\mathbf{b}_{\textrm{R}}$ at the receiver and $i_{\textrm{A}}^{\textrm{R}}$ is the current at the antenna port. $i_{\textrm{A}}^{\textrm{R}}$ is an optimization variable to ensure that the constraint (13) can be satisfied for any given $\mathbf{b}_{\textrm{R}}$.
To handle the constraint (13), we normalize the radiation pattern $\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ and therefore equivalently transform the problem (12)-(18) as
$\underset{\mathbf{e}_{\textrm{R}},\mathbf{b}_{\textrm{R}},\bar{\mathbf{i}}_{\textrm{R}}}{\text{max}\ \ }$ $\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}$ (19)
s.t. $\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\frac{\mathbf{E}_{\textrm{oc}}\bar{\mathbf{i}}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)}{\left\|\mathbf{E}_{\textrm{oc}}\bar{\mathbf{i}}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)\right\|},$ (20)
$\bar{\mathbf{i}}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\left[\begin{array}[]{c}1\\ -\left(\mathbf{Z}_{\textrm{PP}}+\mathbf{Z}_{\textrm{L}}\left(\mathbf{b}_{\textrm{R}}\right)\right)^{-1}\mathbf{z}_{\textrm{PA}}\end{array}\right],$ (23)
$\left[\mathbf{b}_{\textrm{R}}\right]_{q}\in\left\{0,1\right\},\forall q,$ (24)
where the variable $i_{\textrm{A}}^{\textrm{R}}$ is removed because it does not affect the normalized radiation pattern. Substituting (23) into (20), we can find the expression of $\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ as a function of $\mathbf{b}_{\textrm{R}}$ and thus equivalently simplify the problem (19)-(24) as
$\underset{\mathbf{b}_{\textrm{R}}}{\text{max}\ \ }$ $\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}$ (25)
@@
Various algorithms have been proposed and utilized for pixel antenna design and optimization such as the genetic algorithm [7], mixed integer linear programming [38], N-port characteristic mode analysis [39], [40], perturbation sensitivity analysis [41], the adjoint method combined with the method of moving asymptote [42], and reinforcement learning [43]. In this work, to solve the problem (25)-(26), we use an efficient algorithm called the Successive Exhaustive Boolean Optimization (SEBO) [44], which has two steps: 1) cyclically optimizing each block of the binary variables by exhaustive search until convergence, and 2) randomly flipping bits in the converged binary solution to check if there is any other local optimum with a better objective value. The computational complexity of SEBO is $\mathcal{O}\left(N_{\textrm{e}}2^{J}\right)$ where $J$ is the block size and $N_{\textrm{e}}$ is the number of iterations. Increasing $J$ can enhance the optimization performance. Particularly, when $J=Q$, the SEBO becomes the exhaustive search and global optimal solution can be found. However, increasing $J$ will exponentially increase the computational complexity. Thus, there is a trade-off between the computational complexity and optimization performance when using the SEBO to optimize antenna coding. More details on SEBO can be found in [44].
@@ III-C Codebook Design for Antenna Coding
$\mathbf{b}_{\textrm{R}}^{\star}=\underset{\mathbf{b}_{\textrm{R}}\in\mathcal{B}}{\text{argmax}}\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (27)
@@
$\underset{\mathbf{b}_{m}}{\text{max}}\ \$$\mathbb{E}\left[\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}\right]$ (28)
@@
The codebook design problem (28) and (29) is related to the vector quantization problem [45]. To handle the ergodic channel gain in (28), we leverage the sample average approximation method [46]. Specifically, we consider a training set containing $L$ realizations for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ which is a random matrix with each entry following i.i.d. complex Gaussian distribution, given by
@@
where $\mathbf{H}_{\textrm{V}}^{\left[l\right]}$ denotes the $l$th virtual channel realization. Accordingly, we can approximate the expectation of channel gain in (28) as the sample average of channel gain, written as
$\mathbb{E}\left[\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\right|^{2}\right]\approx\frac{1}{L}\sum_{l=1}^{L}\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}^{\star}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}.$ (31)
@@
$\mathcal{H}_{m}$ $=\left\{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mid\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}\right$.
$\left.\qquad\hfill\geq\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m^{\prime}}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall l,\forall m^{\prime}\neq m\right\},\forall m.$ (32)
@@
$\underset{\mathbf{b}_{m}}{\text{max}}\ \$$\sum_{m=1}^{M}\sum_{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\in\mathcal{H}_{m}}\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}$ (33)
@@
Note that to ensure that the sample average accurately approximates the ergodic channel gain, the entries of training set should be the independent realizations of random $\mathbf{H}_{\textrm{V}}$ and the size of training set $L$ should be large enough compared to the codebook size. Otherwise, such approximation is inaccurate and will degrade the codebook design.